C) Type ß error D) Type I error 43) 43) The dean of a major university claims that the mean number of hours students study at her Unlversity (per day) is at most 3.8 hours. If a hypothesis test is performed, how should you Interpret a decision that fails to reject the null hypothesis? A) There is not sufficient evidence to support the claim u s3.8. B) There is not sufficient evidence to reject the claim ? s38. C) There is sufficient evidence to reject the claimus3.8. D) There is sufficient evidence to support the claim u s3.8. 4) 44) The mean age of professors at a university is greater than 51.2 years. If a hypothesis test is performed, how should you interpret a decision that rejects the null hypothesis? A) There is sufficient evidence to support the claim u> 51.2 B) There is not sufficient evidence to reject the claim ? > 512. C) There is sufficient evidence to reject the claim H> 51.2. D) There is not sufficient evidence to support the claim ?> 51.2. 45) 45) A survey of 100 fatal accidents showed that in 24 cases the driver at fault was inadequately insured. Find a point estimate for p, the population proportion of accidents where the driver at fault was inadequately insured A) 0.24 B) 0.316 C) 0.194 D) 0.76 Provide an appropriate response. Round the test statistic to the nearest thousandth. 46) 46) Compute the standardized test statistic, X2, to test the claim ?, 11.4 ?fn-18, S2-162, and a 0.01. C) 33.233 D) 28.175 A) 43.156 B) 24.158 12, s. 18, and 47) 47) Compute the standardized test statistic, X2, to test the claim ?2-215 in a 0.05 D) 9.209 C) 18.490 A) 0.492 B) 12.961
media%2F0ac%2F0acd311d-0875-45ad-bcfe-4f
Answer #1

Ans:

43)Option B is correct.

There is not sufficient evidence to reject the claim \mu \leq 3.8

44)Option A is correct.

There is sufficient evidence to support the claim \mu > 51.2

45)

Point estimate=sample proportion=24/100=0.24

Option A is correct.

46)

Test statistic:

Chi square score=(18-1)*16.2/11.4=24.158

Option B is correct.

47)

Test statistic:

Chi square score=(12-1)*18/21.5=9.209

Option D is correct.